OpenStudy (anonymous):

plzzz helpppp! Choose the equation of the line passing through the point (1, -5) and perpendicular to y = 1/5x - 10. y = 5x + 10 y = 5x y = -5x - 10 y = -5x

OpenStudy (anonymous):

if the line is to be perpendicular to the line with slope of 1/5, what must the slope be then? do you know the relationship between perpendicular slopes?

OpenStudy (anonymous):

slope is 5

OpenStudy (anonymous):

no.... the relationship is that the product of perpendicular slopes is -1. in other words, \[\huge m_1 \cdot m_2=-1 \rightarrow \frac{1}{5} \cdot m_2=-1 \] so what must \(\large m_2= \) ????